OpenStudy (anonymous):

suppose S is between R and T. Use the segment addition postulate to solve for each variable. RS= 7y-4 , ST= y+5 , RT= 28

jimthompson5910 (jim_thompson5910):

using the segment addition postulate, we can say RS + ST = RT (7y-4) + (y+5) = 28 7y - 4 + y + 5 = 28 8y + 1 = 28 I'll let you finish